A-ONE publisher and sells books. The management is in the process of printing a new book to be marked. 

The book may be bound by either cloth or paperback. Each cloth bound book sold out bring a profit of Rs 

60 and each paper bound book contributes Rs 35. It takes 10 minutes to bind a cloth cover and 8 minutes 

to bind a paperback. The total available time for binding is 20 hours. After considerable market surrey, it is 

predicated that the cloth cover sales will not exceed 800 copies and the paperback sales will be no more 

than 1200 copies. Formulate as an LP model with an appropriate objective function and constraints.


1
Expert's answer
2021-12-07T13:08:47-0500

LP model: maximize profit:

max "60x+35y"

subject to

"10x+8y\le 20\cdot60=1200" , total available time for binding

"x\le 800"

"y\le 1200"


x is number of cloth bound books

y is number of paper bound books